OpenStudy (anonymous):

Solve the following exponential equations 1) 10^2x + 27 =190 2) 4e^x-2 +5 =70

OpenStudy (anonymous):

You are solving for x right? Isolate x onto one side. Try to do that first and come back when you get stuck or confused about a step.

OpenStudy (campbell_st):

1. subtract 27 from both sides \[10^{2x} = 163\] take the base 10 log of both sides use the fact \[\log_{10}(10^{2x}) = 2x\] then \[2x = \log _{10}163\] \[x = \frac{\log _{10}(163)}{2}\]

OpenStudy (campbell_st):

Question 2 same process do all the arithmetic 1st take the base e log to get an answer

OpenStudy (anonymous):

so, is the final answer of the seconed equation ganna be x= ln4 (65)/-2 ?!

OpenStudy (campbell_st):

\[4e^{x-2} = 65\] then divide by 4 \[e^{x -2} = \frac{65}{4}\] \[\ln(e^{x - 2}) = x-2\] then \[x - 2 = \ln(\frac{65}{4})\] \[x = 2 + \ln(\frac{65}{4})\]